### v2

# Bayesian Inference General Procedures for A Single-subject Test Study

Abstract
Abnormality detection in the identification of a single-subject which deviates from the majority of the dataset that comes from a control group is a critical problem. A common approach is to assume that the control group can be characterised in terms of standard Normal statistics and the detection of single abnormal subject is in that context. But in many situations the control group can not be described in terms of Gaussian statistics and the use of standard statistics is inappropriate. This paper presents a Bayesian Inference General Procedures for A Single-Subject Test (BIGPAST), designed to mitigate the effects of skewness under the assumption that the dataset of control group comes from the skewed Student’s $t$ distribution. BIGPAST operates under the null hypothesis that the single-subject follows the same distribution as the control group. We assess BIGPAST’s performance against other methods through a series of simulation studies. The results demonstrate that BIGPAST is robust against deviations from normality and outperforms the existing approaches in terms of accuracy. This is because BIGPAST can effectively reduce model misspecification errors under the skewed Student’s $t$ assumption. We apply BIGPAST to a MEG dataset consisting of an individual with mild traumatic brain injury and an age and gender-matched control group, demonstrating its effectiveness in detecting abnormalities in the single-subject.
Keywords— Bayesian inference, Skewed Student’s $t$ distribution, Single-subject test, Control group, Magnetoencephalography (MEG), Jeffery’s prior, and False discovery rate
1 Introduction
Abnormality detection is a fundamental task in many scientific fields, including medicine, psychology, and neuroscience (Scholz and Stephens, 1987; Crawford and Howell, 1998; Crawford and Garthwaite, 2007). One example of such abnormality detection lies in the detection of traumatic brain injury (TBI), which is a particularly common problem, with over 2.8 million individuals a year in America presenting at an Emergency room (Taylor et al., 2017). We exemplify the use of skewed Student’s $t$ statistics to the study of brain activity as a potential biomarker for the identification of a brain injury. We use magnetoencephalography (MEG) as the technology for measuring brain activity and compare single individual to normal controls from age and gender-matched cohorts. Existing literature normally assumes that the dataset from the control group follows a Normal distribution. However, extensive practical evidence shows that the naive Normal distribution assumption is often violated. This violation is frequently observed in brain Magnetoencephalography (MEG) datasets. Therefore, the current approaches using Normal statistics could not be directly applied. Unlike the existing literature, we assume that the dataset from the control group belongs to a more general distributional family - the skewed Student’s $t$ distribution, which includes normal distribution and Student’s $t$ as a special case. The benefit of this skewed Student’s $t$ distributional assumption is that it could cover a variety of empirical behaviours and is thus more suitable for practical needs. This generalisation is essential since many brain Magnetoencephalography (MEG) datasets show a significant level of asymmetry in the observation distribution and hence are highly skewed. It is also the case that the high skewness not only appears in a MEG dataset, but also in other fields, such as cybersecurity (Kim and Song, 2023), medicine (Lu and Huang, 2017), and neuroscience (Buzsáki and Mizuseki, 2014) to name only a few.
Under the assumption of the skewed Student’s $t$ distribution in this article, we describe the statistical issues involved in the comparison of a single individual with a large normal control group. The statistical issues arise from the interplay between observation from single-subject(whether well-mixed or not), the type of hypothesis (two-sided, less, greater), and the sign of skewness parameter which we will discuss one-by-one in the following paragraphs.
Under the assumption that dataset observed from the control group follows a Normal distribution, there exist several methods that are able to compare a single-subject against a control group. Crawford and Howell (1998) proposed a $t$-score test defined as follows:
$t=\frac{x^{*}-\bar{x}}{s/\sqrt{(n+1)/n}},$
where $x^{*}$ represents the observation of the single-subject, $\bar{x}$ denotes the average of $n$ observations from the control group, and $s$ stands for the standard deviation of the dataset from the control group. The test score $t$ follows a Student’s $t$ distribution with $n-1$ degrees of freedom. Let $\alpha$ be the prespecified significance level, define the $p$ value as $\mathbb{P}(t>|t_{1-\alpha/2}(n-1)|)$ where $t_{1-\alpha/2}(n-1)$ is the $(1-\alpha/2)$-th critical value of Student’s $t$ distribution with $n-1$ degrees of freedom. If the $p$ value is less than $\alpha$, then one can claim that $x^{*}$ has an abnormality. Moreover, under the assumption of skewed Student’s $t$ distribution, Crawford et al. (2006) examined the effects of departures from normality on testing for a deficit in a single-subject via using a Leptokurtic distribution with different skewness parameters. They asserted that the $t$ score (Crawford and Howell, 1998) outperforms the $z$ score in terms of Type I error, and the $t$ score shows robustness as good as the result under the Normal assumption. Additionally, Crawford and Garthwaite (2006) investigated the statistical power of testing for an abnormality in a single-subject using Monte Carlo simulations. They claimed that the power of $t$ score performs slightly worse compared with the results under Normal assumption.
The simulation results for Type I error in Crawford et al. (2006) and power in Crawford and Garthwaite (2006) led to the assertion that the $t$ score (Crawford and Howell, 1998) is robust even when the underlying distribution is skewed Student’s $t$ distribution. However, the simulation settings in Crawford et al. (2006)Crawford and Garthwaite (2006) are not comprehensive (ground true observations of a single subject are not well-mixed) to evaluate the performance of $t$ score under skewed Student’s $t$ distribution.
Predicated
Positive Negative
Actual Positive NA NA
Negative $n_{01}$ $n_{00}$
(a) Type I error evaluation in Crawford et al. (2006)
Predicated
Positive Negative
Actual Positive $n_{11}$ $n_{10}$
Negative NA NA
(b) Power evaluation in Crawford and Garthwaite (2006)
Table 1: The data generation mechanism of single-subject observations in Crawford et al. (2006)Crawford and Garthwaite (2006) respectively. ‘NA’ means that the observations in the cell is not available. ‘Positive’ means that the single-subject is claimed as abnormal, while ‘Negative’ means that the single-subject is claimed as normal.
For instance, in evaluating Type I error, Crawford et al. (2006) generates all single-subject observations from the distribution of the control group, see Table 1(a). In  Table 1(a), $n_{01}$ is the number of false positive (FP) while $n_{00}$ is the number of true negative (TN). The false positive rate (FPR, i.e., Type I error) is defined as $n_{01}/(n_{01}+n_{00})$. As there is no actual positive observations in Table 1(a), the true positive rate (TPR, i.e., power) is undefined for the simulation settings in Crawford et al. (2006). Therefore, the comparison of Type I error rates in Crawford et al. (2006) is not thoroughly discussed, as the study did not control for statistical power. Conversely, for the power analysis in Crawford and Garthwaite (2006), all single-subject observations are generated from a distribution that is different from the distribution of control group, see Table 1(b). In Table 1(b), $n_{11}$ is the number of true positive (TP) while $n_{01}$ is the number of false negative (FN). The true positive rate (TPR, i.e., power) is defined as $n_{11}/(n_{11}+n_{10})$. However, the Type I error is undefined. Again, Crawford and Garthwaite (2006) did not fully discuss the comparison of the power evaluation due to the absence of Type I error. Neither Crawford et al. (2006) nor Crawford and Garthwaite (2006) discussed that the ground true observations of a single-subject are well-mixed. For instance, if 50% of the single-subject observations are generated from underlying distribution of the control group data, the other 50% of observation comes from a different distribution, the Type I error and power of the $t$ score test would differ significantly. A direct evidence illustrated in Figure 2(c) shows that the powers (in median) of Crawford-Garthwaite (CG) method and BIGPAST under two-sided test are around 0.62 and 0.96 respectively. To address this issue, all the simulation settings (unless specified) of the single-subject in this paper are well-mixed, i.e., the single-subject observations are generated from a mixture of the null and alternative distribution. The well-mixed setting is more realistic and can provide a comprehensive evaluation of the performance of BIGPAST test under skewed Student’s $t$ assumption.
Before delving into the rationale behind the assumption of the skewed Student’s $t$ distribution, it is essential to review the definitions of heavy tails and light tails within the context of the skewed Student’s $t$ distribution. Let $F(x)$ be the cumulative distribution function of random variable $X$. The right tail distribution is denoted as $\overline{F}(x)\coloneqq\Pr(X>x)$ while the left tail distribution is denoted as $\underline{F}(x)\coloneqq\Pr(X\leq x)$. $F$ is said to have a right heavy tail if $\lim_{x\to+\infty}\exp({tx})\overline{F}(x)=\infty$ for all $t>0$; have a right light tail if $\exp({tx})\overline{F}(x)=O(1)$ for large positive $x$ and some constant $t>0$. Similarly, $F$ is said to have a left heavy tail if $\lim_{x\to-\infty}\exp({tx})\underline{F}(x)=\infty$ for all $t>0$; have a left light tail if $\exp({tx})\underline{F}(x)=O(1)$ for small negative $x$ and some constant $t>0$. In other words, light tails decay to zero much faster than the exponential distribution, whereas heavy tails decay much slower, see Figure 1.
(a) Positive skewness: $\alpha=2$ with $\nu=3$
(b) Negative skewness: $\alpha=-2$ with $\nu=5$
Figure 1: The heavy and light tails of the standard skewed Student’s $t$ distribution for both positive and negative skewness parameters. The orange lines represent the standard Student’s $t$ density function, the green lines represent the curve of $\exp({-{\left\lvert x\right\rvert}})$. The vertical lines are quantiles: $t_{0.05},t_{0.95}$ (resp. $t_{0.05}^{*},t_{0.95}^{*}$) are 5% and 95% quantiles of the standard Student’s $t$ distribution (resp. the standard skewed Student’s $t$ distribution). In (a), $t_{0.05}=-2.353,t_{0.95}=2.353$, $t_{0.05}^{*}=-0.398,t_{0.95}^{*}=3.169$. In (b), $t_{0.05}=-2.015,t_{0.95}=2.015$, $t_{0.05}^{*}=-2.568,t_{0.95}^{*}=0.370$.
It is well-known that the Student’s $t$ distribution is symmetric and has two heavy tails at both sides. The skewed Student’s $t$ distribution is a generalisation of the Student’s $t$ distribution (Azzalini and Capitanio, 2014). The skewness parameter $\alpha$ in Equation 2 changes the symmetry of Student’s $t$ distribution given the degree of freedom $\nu$. When $\alpha=0$ and $\nu$ is finite, the skewed Student’s $t$ distribution reduces to Student’s $t$ distribution which has both heavy tails; When $\alpha=0$ and $\nu=+\infty$, the skewed Student’s $t$ distribution reduces to normal distribution which has both lighter tails; When $|\alpha|=\infty$ and $\nu<+\infty$, the skewed Student’s $t$ distribution reduces to half Student’s $t$ distribution which can be tuned to allow for a heavy tail ($\nu\to 1$) or a light tail (large $\nu$). Therefore, for other settings of ($\alpha,\nu$), the skewed Student’s $t$ distribution may have one light tail and one heavy tail. The sign of $\alpha$ dictates whether the heavy tail is on the left or right side of the distribution, see Figure 1. In Figure 1(a), the skewness is positive, the standard skewed Student’s $t$ distribution has a right heavy tail and a left light tail. Conversely, in Figure 1(b), the skewness is negative, the standard skewed Student’s $t$ distribution has a right light tail and a left heavy tail, see the zoomed insets in Figures 1(a) and 1(b) respectively.
In this paper, we consider three kinds of alternative hypothesis test: “less than” ($<$), “greater than” ($>$) and “two-sided”. For clarity, we refer to the alternative hypothesis with symbol ($<$) as the one-sided test (less), and the alternative hypothesis with symbol ($>$) as the one-sided test (greater). Under the assumption of symmetric distributions (e.g. Normal or Student’s $t$), one-sided test, either “less” or “greater”, share the same absolute critical value given the same significance level. However, under the assumption of skewed Student’s $t$ distribution, the absolute critical value of the one-sided test (less) is different from the absolute critical value of the one-sided test (greater) due to relationship between skewness and heavy tail. Hence, under the assumption of a skewed Student’s $t$ distribution, there are six combinations of hypothesis tests: the one-sided test (less), the one-sided test (greater), and the two-sided test, each with either a positive or negative skewness parameter.
Now, let us look into the parameter settings in Crawford et al. (2006). They set $\alpha<0$ and only considered the one-sided test (less). Both the observations of control group and single-subject are generated from the same skewed Student’s $t$ distribution with the same skewness parameter. Furthermore, the observations of single-subject are all negative as described in Table 1(a). They used the $t$ statistic to test the abnormalities of the single-subject. Figure 1(b) describes a similar skewness parameter setting as in Crawford et al. (2006). Given the significance level 0.95, the critical value of the one-sided test (less) for $t$ test is $t_{0.05}=-2.015$ while the critical value of the one-sided test (less) for skewed $t$ test is $t_{0.05}^{*}=-2.568$. Apparently, $t_{0.05}>t^{*}_{0.05}$, which implies that without considering the randomness, if some single-subject observation $x^{*}$ is claimed as abnormal by $t_{0.05}^{*}$, then $x^{*}$ must be claimed as abnormal by $t_{0.05}$. Let’s define the model mis-specification error for one-sided test (less) as the probability $\Pr_{t}(t_{0.05}^{*}\leq x\leq t_{0.05})$ where $\Pr_{t}(\cdot)$ is the probability under the Student’s $t$ distribution. In Figure 1(b), the model mis-specification error for one-sided test (less) is $\Pr_{t}(-2.568\leq x\leq-2.015)\approx 0.0249$. However, if we consider the one-sided test (greater) under the same settings, the model mis-specification error for one-sided test (greater) is $\Pr_{t}(0.0370\leq x\leq 2.015)\approx 0.3131$. The model mis-specification error for the one-sided test (greater) is almost 12 times higher than that for the one-sided test (less). This means that the one-sided test (less) is more robust than the one-sided test (greater) when $\alpha<0$. Under the settings of $\alpha<0$, one-sided test (less) and not well-mixed observations of single-subject, Crawford et al. (2006) claimed that the $t$ score is robust in terms of Type I error even under skewed Student’s $t$ distribution. However, this assertion does not hold if the one-sided test (less) is replaced with the one-sided test (greater). For the power comparison, Crawford and Garthwaite (2006) make the similar statement as well. However, by using the similar discussion above, we can find that the robustness statement for the power only holds for the specific settings in their paper.
Moreover, we investigate the model misspecification error between normal and skewed Student’s $t$ distributions as well in Figure 3. The regions shaded in light-red colour in Figures 3(a) and 3(b) represent the model misspecification errors that occur when the skewed Student’s $t$ distribution is assumed, but the $z$ score test is used. Apparently, the $z$ score will always have model misspecification error under the skewed Student’s $t$ distribution assumption. In fact, neither the $z$ score test, the $t$ score test, nor even the Crawford-Garthwaite Bayesian method (Crawford and Garthwaite, 2007) can eliminate the model misspecification error.
The rationale for this paper’s assumption of the skewed Student’s $t$ distribution can be attributed to the fact that it can help eliminate model misspecification errors, provided that the underlying distribution is a special case of the skewed Student’s $t$ distribution. These special cases encompass the regular Student’s $t$ distribution ($\alpha=0$), skewed normal distribution ($\nu=\infty$), half normal distribution ($\alpha=\pm\infty,\nu=\infty$), and normal distribution ($\alpha=0,\nu=\infty$) as special cases. To show the elimination of model misspecification error, simulation results with normal assumption but using BIGPAST method can be found in Table 9.
We can conclude that the interplay between skewness (whether positive or negative), the type of hypothesis (‘two-sided’, ‘less’, or ‘greater’), and the single-subject observations (whether well-mixed or not) significantly influences both the Type I error rate and the power of the $t$ score test under the assumption of skewed Student’s $t$ distribution. To investigate the influence, this paper introduces a Bayesian Inference General Procedures for A Single-subject Test (BIGPAST) framework under the skewed Student’s $t$ assumption to fully discuss the test performance in different settings (i.e., the combinations of the sign of $\alpha$, type of hypothesis and well-mixed single-subject observations). The BIGPAST framework is analogous to the Bayesian inference framework proposed by Crawford and Garthwaite (Crawford and Garthwaite, 2007), with the key distinction being our assumption of the skewed Student’s $t$ distribution. Additionally, we propose a Jefferys prior which is defined by explicit mathematical formulas. We evaluate BIGPAST’s performance against the existing approaches through simulation studies. The results demonstrate that BIGPAST is robust against departures from normality and surpasses the existing methods based on the assumption of a normal distribution in terms of Type I error, power and accuracy.
The remainder of the paper is organised as follows. In Section 2, we introduce the Bayesian inference procedures for comparing a single-subject with a control group under the skewed Student’s $t$ distribution. In Section 3, we conduct a comprehensive set of simulation studies to evaluate the performance of the proposed BIGPAST method. We then apply the BIGPAST method to a MEG dataset comprising an individual with mild traumatic brain injury and a control group in Section 4. Finally, we summarise our findings and discuss future research directions in Section 5.
2 Bayesian Inference Procedures for Single-subject’s Abnormality Detection
In this section, we introduce a Bayesian inference procedures to compare a single-subject with a control group, under the assumption that the control group observations follow a skewed Student’s $t$ distribution. First, we should verify whether the assumption holds. For practical applications, we recommend employing a modified version of the general goodness-of-fit test (Chen and Balakrishnan, 1995) tailored explicitly for the skewed Student’s $t$ distribution in the Appendix A. From now on, we assume that the control group data follows a skewed Student’s $t$ distribution in the subsequent sections.
2.1 Bayesian Inference General Procedures for a Single-subject Test (BIGPAST)
Let $x^{*}$ denote the observation of a single-subject. Our goal is to test whether $x^{*}$ is a random sample from the same distribution as the control group. Let $f_{c}(\alpha,\nu,\xi,\omega)$ represent the skewed Student’s $t$ density function of $x_{1},x_{2},\ldots,x_{n}$. The hypothesis testing can be formulated as follows:
$\mathcal{H}_{0}:x^{*}\sim f_{c}(\alpha,\nu,\xi,\omega)\quad v.s.\quad\mathcal{H}_{1}:x^{*}\not\sim f_{c}(\alpha,\nu,\xi,\omega),$
where $\not\sim$ represents that $x^{*}$ is not a random sample from $f_{c}(\alpha,\nu,\xi,\omega)$. We propose the Bayesian inference general procedures for a single-subject test under the null hypothesis $\mathcal{H}_{0}$. The procedures are as follows:
  1. 1.
    Choose the proposed Jeffery prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ in Section 2.2 as the joint prior of $\alpha,\nu,\xi,\omega$.
  2. 2.
    Compute the posterior distribution of $\alpha,\nu,\xi,\omega$ based on the control group observations $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ and the prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$. By Bayes’s Theorem, we have
    $\boldsymbol{\pi}(\alpha,\nu,\xi,\omega|\mathbf{x})\propto l(\mathbf{x}|\alpha,\nu,\xi,\omega)\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega),$
    where $l(\mathbf{x}|\alpha,\nu,\xi,\omega)\coloneqq\prod_{i=1}^{n}f_{c}(x_{i}|\alpha,\nu,\xi,\omega)$ is the likelihood function.
  3. 3.
    Draw $m$ samples of $\alpha,\nu,\xi,\omega$ from the posterior distribution $\boldsymbol{\pi}(\alpha,\nu,\xi,\omega|\mathbf{x})$ by Metropolis-Hastings algorithm (Hastings, 1970) with normal proposal, denoted as $\alpha_{j},\nu_{j},\xi_{j},\omega_{j}$, $j=1,2,\ldots,m$.
  4. 4.
    For each $\alpha_{j},\nu_{j},\xi_{j},\omega_{j}$, draw $s$ random samples from the skewed Student’s $t$ distribution, denoted as $x_{1j},x_{2j},\ldots,x_{sj}$, $j=1,2,\ldots,m$. Sort the observations $x_{1j},x_{2j},\ldots,x_{sj}$, $j=1,2,\ldots,m$, in ascending order and label them as $x_{(1)},x_{(2)},\ldots,x_{(B)}$ where $B=ms$.
  5. 5.
[…]
Remarks
When $\alpha\to+\infty$ or $\infty$- ∞, the skewed Student’s $t$ distribution degenerates to half Student’s $t$ distribution. Then, the two-sided credible interval in Step 5 above should be $[x_{(1)},x_{(\lceil B(1-\alpha^{*})\rceil)}]$ if $\alpha=+\infty$ or $[x_{(\lfloor B\alpha^{*}\rfloor)},x_{(B)}]$ if $\alpha=-\infty$. Furthermore, procedures 1-3 above can be extended to other distributions, provided that a prior can be assigned for a specific density distribution.
2.2 Jefferys Priors
In Bayesian probability, the Jeffreys prior (Jeffreys, 1946) refers to a prior density function that is proportional to the square root of the determinant of the Fisher information matrix. In this section, we proposed an explicit Jeffery prior for the parameters of skewed Student’s $t$ distribution and compare it to the existing priors.
Let $f(z|\alpha,\nu)$ be the non-linear skewed Student’s $t$ density function (Azzalini and Capitanio, 2003), defined as follows:
$f(z|\alpha,\nu)\coloneqq 2t(z|\nu)T(w|\nu+1),\quad z\in\mathbb{R},\alpha\in\mathbb{R},\nu>0,$ (1)
where $r(z,\nu)=\sqrt{(\nu+1)/(\nu+z^{2})}$ and $w=\alpha zr(z,\nu)$, $t(\cdot|\nu)$ represents the density function of $t$ distribution with degree of freedom $\nu$, and $T(\cdot|\nu+1)$ represents the cumulative density function of $t$ distribution with degree of freedom $\nu+1$. Let $\xi$ and $\omega$ be the location and scale parameters, with the substitution $z=(x-\xi)/\omega$ in Equation 1, then the skewed Student’s $t$ density function is given by
$f(x|\alpha,\nu,\xi,\omega)\coloneqq\frac{1}{\omega}f\left(\frac{x-\xi}{\omega}|\alpha,\nu\right).$ (2)
In the literature, two types of priors are commonly used for $\xi$ and $\omega$: the partial information prior (Sun and Berger, 1998) and the independence Jefferys prior (Rubio and Steel, 2014). Both of these priors assume that the joint prior of $\xi$ and $\omega$, denoted as $\boldsymbol{\pi}(\xi,\omega)$, is proportional to $\omega^{-1}$. Consequently, the joint prior of $\alpha,\nu,\xi,\omega$ can be expressed as follows:
$\boldsymbol{\pi}(\alpha,\nu,\xi,\omega)\propto\omega^{-1}\boldsymbol{\pi}(\alpha,\nu).$
The joint prior of $\alpha$ and $\nu$, $\boldsymbol{\pi}(\alpha,\nu)$, may be specified by the Jefferys prior (Branco et al., 2013) or the natural informative prior (Dette et al., 2018). Branco et al. (2013) proposed the following prior for general non-linear skewed Student’s $t$ density function:
$\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)\propto\omega^{-1}\left(8\alpha^{2}+\pi^{2}\right)^{-3/4}\boldsymbol{\pi}(\nu),$
where $\boldsymbol{\pi}(\nu)$ is proposed by Fonseca et al. (2008) for the Student’s $t$ distribution and given by
$\boldsymbol{\pi}(\nu)\propto\left(\frac{\nu}{\nu+3}\right)^{1/2}\left\{\psi^{(1)}\left(\frac{\nu}{2}\right)-\psi^{(1)}\left(\frac{\nu+1}{2}\right)-\frac{2(\nu+3)}{\nu(\nu+1)^{2}}\right\}^{1/2},$
where $\psi^{(1)}(\cdot)$ is the trigamma function, i.e., the second derivative of $\log(\Gamma(\cdot))$. Here, $\Gamma(\cdot)$ is the Gamma function. When $\alpha\rightarrow+\infty$, the order of $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$ with respect to $\alpha$ is $O(\alpha^{-3/2})$ (Branco et al., 2013).
Dette et al. (2018) proposed the so-called natural informative prior of the parameter $\alpha$ based on the total variance distance, which is given by
$BTV(\alpha|\theta_{1},\theta_{2})=\frac{1}{B(\theta_{1},\theta_{2})}\left(\frac{\tan^{-1}(\alpha)}{\pi}+\frac{1}{2}\right)^{\theta_{1}-1}\left(\frac{-\tan^{-1}(\alpha)}{\pi}+\frac{1}{2}\right)^{\theta_{2}-1}\frac{1}{\pi(1+\alpha^{2})},$
where $B(\cdot,\cdot)$ represents the Beta function. When $\theta_{1}=1,\theta_{2}=1$, the prior $BTV(\alpha|1,1)$ reduces to the standard Cauchy prior. Similarly, Dette et al. (2018) constructed the joint prior of $\alpha,\nu,\xi,\omega$ based on $BTV(\alpha|1,1)$ as
$\boldsymbol{\pi}^{D}(\alpha,\nu,\xi,\omega)\propto\omega^{-1}\pi^{-1}\left(1+\alpha^{2}\right)^{-1}\boldsymbol{\pi}(\nu),$
which is order of $O(\alpha^{-2})$ when $\alpha\rightarrow+\infty$.
A notable concern with $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$ and $\boldsymbol{\pi}^{D}(\alpha,\nu,\xi,\omega)$ is that the marginal prior of $\nu$ is independent of $\alpha$. However, given that the skewed Student’s $t$ distribution is a generalisation of the Student’s $t$ distribution, the marginal prior of $\nu$ should ideally depend on $\alpha$. To address this issue, we propose the Jefferys prior for the skewed Student’s $t$ density function, as detailed in Theorem 2.1.
Theorem 2.1.
Let $f(z|\alpha,\nu)$ be the non-linear skewed Student’s $t$ density function, defined as follows:
$f(z|\alpha,\nu)\coloneqq 2t(z|\nu)T(w|\nu+1),\quad z\in\mathbb{R},\alpha\in\mathbb{R},\nu>0,$
where $r(z,\nu)=\sqrt{(\nu+1)/(\nu+z^{2})}$ and $w=\alpha zr(z,\nu)$. Given any $\nu>0$, let $\boldsymbol{\pi}^{J}(\alpha,\nu)$ represent the joint Jefferys prior for $\alpha$ and $\nu$, we have
$\boldsymbol{\pi}^{J}(\alpha,\nu)\propto\sqrt{I_{\alpha\alpha}I_{\nu\nu}-I_{\alpha\nu}^{2}},$ (3)
and the joint prior of $\alpha,\nu,\xi,\omega$ is given by
$\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)\propto\omega^{-1}\boldsymbol{\pi}^{J}(\alpha,\nu),$ (4)
where
$I_{\nu\nu}$ $=\mathbb{E}\left[h^{2}(w)\frac{\alpha^{2}z^{2}(\nu+1)}{4\left(\nu+z^{2}\right)^{3}}\right]+\frac{1}{4}\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)^{2}\right]+\frac{1}{4}\mathbb{E}\left[\left(\frac{z^{2}-1}{\nu+z^{2}}\right)^{2}\right]+d^{2}_{\nu}$
$\qquad+\frac{1}{2}\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)\left(\frac{z^{2}-1}{\nu+z^{2}}\right)\right]+d_{\nu}\mathbb{E}\left[\log\frac{\nu}{\nu+z^{2}}\right]$
$I_{\alpha\nu}$ $=-\frac{\pi\alpha\Gamma\left(\frac{\nu}{2}+1\right)^{2}\left((\nu+4)H_{5}-\frac{\left(\alpha^{2}(2\nu+3)+(\nu+3)\sigma_{\nu+1}^{2}\right)H_{6}}{\sigma_{\nu+1}^{2}}\right)}{8\left(\alpha^{2}+\sigma_{\nu+1}^{2}\right)\Gamma\left(\frac{\nu+1}{2}\right)\Gamma\left(\frac{\nu+5}{2}\right)},$
and
$\mathbb{E}\left[\log\frac{\nu}{\nu+z^{2}}\right]$ $=\psi\left(\frac{\nu}{2}\right)-\psi\left(\frac{\nu+1}{2}\right)$
$\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)\left(\frac{z^{2}-1}{\nu+z^{2}}\right)\right]$ $=-\frac{1}{\nu^{2}+\nu}$
$\mathbb{E}\left[\left(\frac{z^{2}-1}{\nu+z^{2}}\right)^{2}\right]$ $=\frac{1}{\nu^{2}+3\nu}$
$\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)^{2}\right]$ $=\left(\psi\left(\frac{\nu}{2}\right)-\psi\left(\frac{\nu+1}{2}\right)\right)^{2}+\psi^{(1)}\left(\frac{\nu}{2}\right)-\psi^{(1)}\left(\frac{\nu+1}{2}\right)$
$\mathbb{E}\left[h^{2}(w)\frac{\alpha^{2}z^{2}(\nu+1)}{4\left(\nu+z^{2}\right)^{3}}\right]$ $=\frac{\pi^{5/2}\Gamma\left(\frac{\nu}{2}+1\right)\left((\nu+3)H_{1}-(\nu+3)H_{2}-H_{3}+H_{4}\right)}{8\nu^{2}\Gamma\left(\frac{\nu+5}{2}\right)B\left(\frac{\nu}{2},\frac{1}{2}\right)B\left(\frac{\nu+1}{2},\frac{1}{2}\right)^{2}}.$
Furthermore, we define $d_{\nu}\coloneqq-\frac{1}{2}\psi\left(\frac{\nu}{2}\right)+\frac{1}{2}\psi\left(\frac{\nu+2}{2}\right)+2c_{\nu}*g_{1}(\nu)$, where $c_{\nu}$ is a constant that depends solely on $\nu$. The hypergeometric functions, denoted as $H_{1},H_{2},H_{3},H_{4},H_{5},H_{6}$, along with $h^{2}(w)$ and $g_{1}(\nu)$ are defined in the proof. The term $\sigma^{2}_{\nu+1}$ is defined in Lemma A.1. The functions $\psi(\cdot)$ and $\psi^{(1)}(\cdot)$ represent the digamma and trigamma function respectively.
The proof of Theorem 2.1 can be found in Appendix. Upon applying a Taylor expansion, we find that as $\alpha\rightarrow 0$, the order of $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ with respect to $\alpha$ is $O(1+\alpha^{2})$. Conversely, as $\alpha\rightarrow+\infty$, the order becomes $O(\alpha^{-3/2})$, aligning with the findings of Branco et al. (2013). The computation of $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ relies solely on the basic functions provided by the Python packages scipy and numpy. As such, the formulas in Theorem 2.1 can be directly implemented to define a function for computing Jefferys prior. Additionally, we have developed a Python package, skewt-scipy (https://pypi.org/project/skewt-scipy/), specifically for the skewed Student’s $t$ distribution.
For the Bayesian estimation of $\nu$, we anticipate that $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ will outperform $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$. Because $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ depends on $\alpha$, unlike $\boldsymbol{\pi}(\nu)$ in $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$ is independent of $\alpha$. This is verified by the comparison study in Section 3. The Jefferys prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ is a suitable choice for the skewed Student’s $t$ density function. The BIGPAST based on Jefferys prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ is robust to detect the departures from normality and outperforms the existing approaches based on the assumption of normal distribution.
3 Simulation
3.1 Jefferys Prior Comparison Study
In this section, we conduct a comparative analysis of our proposed prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ against the priors $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$ and $\boldsymbol{\pi}^{D}(\alpha,\nu,\xi,\omega)$. We evaluate these priors under six different parameter settings for $(\alpha,\nu)$: $(-1,1)$, $(-10,10)$, $(-30,30)$, $(-50,50)$, $(-1,50)$, and $(-50,1)$. For each setting, the location and scale parameters are fixed at $\xi=-2,\omega=\sqrt{2}$. To simplify the notation, we use $\boldsymbol{\pi}^{J}$, $\boldsymbol{\pi}^{B}$, and $\boldsymbol{\pi}^{D}$ as shorthand for $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$, $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$, and $\boldsymbol{\pi}^{D}(\alpha,\nu,\xi,\omega)$, respectively. For each parameter setting, we generate $N=1000$ samples, each with a sample size of $n=500$, from the non-linear skewed Student’s $t$ density function. We then numerically minimise the negative logarithm of the posterior distribution to obtain the maximum posterior estimators of $\alpha,\nu,\xi,\omega$. The performance of different priors is evaluated using the mean absolute deviation (MAD) from the true parameter. The MAD is calculated as $1/N\sum_{i=1}^{N}{\lvert\hat{\theta}_{i}-\theta\rvert}$, where $\theta$ is the true parameter and $\hat{\theta}_{i}$ is the $i$-th maximum posterior estimator of $\theta$. Our comparative study also includes the maximum likelihood estimators (MLE). The results are presented in Table 2.
As demonstrated in Table 2, when $|\alpha|$ and $\nu$ are less than 30, the performance of $\boldsymbol{\pi}^{J}$ aligns closely with that of $\boldsymbol{\pi}^{B}$ and $\boldsymbol{\pi}^{D}$ in terms of estimating $\alpha$ and $\nu$. However, for larger values of $|\alpha|$ and $\nu$, $\boldsymbol{\pi}^{J}$ outperforms both $\boldsymbol{\pi}^{B}$ and $\boldsymbol{\pi}^{D}$. Furthermore, $\boldsymbol{\pi}^{J}$, $\boldsymbol{\pi}^{B}$, and $\boldsymbol{\pi}^{D}$ all outperform the maximum likelihood estimators (MLE) for larger $|\alpha|$ and $\nu$. In terms of estimating the location $\xi$ and scale $\omega$, the mean absolute deviations (MADs) of all priors are comparable to the MLE. These results underscore the efficacy of $\boldsymbol{\pi}^{J}$ for the non-linear skewed Student’s $t$ density function, particularly when $|\alpha|$ and $\nu$ exceed 30.
Parameters Priors $(\alpha,\nu)$
$(-1,1)$ $(-10,10)$ $(-30,30)$ $(-50,50)$ $(-1,50)$ $(-50,1)$
$\alpha$ $\boldsymbol{\pi}^{J},c_{\nu}=0$ 0.156 1.989 9.402 11.839 0.471 8.415
$\boldsymbol{\pi}^{J},c_{\nu}=1$ 0.162 2.009 9.105 11.072 0.457 7.703
$\boldsymbol{\pi}^{B}$ 0.142 1.901 9.399 14.267 0.497 11.489
$\boldsymbol{\pi}^{D}$ 0.142 1.913 9.122 14.362 0.532 11.807
Uniform 0.146 2.451 236.974 375.183 0.385 612.510
$\nu$ $\boldsymbol{\pi}^{J},c_{\nu}=0$ 0.079 2.612 6.130 6.656 14.321 0.064
$\boldsymbol{\pi}^{J},c_{\nu}=1$ 0.085 2.418 5.474 6.210 8.961 0.063
$\boldsymbol{\pi}^{B}$ 0.064 2.642 15.143 20.924 35.187 0.061
$\boldsymbol{\pi}^{D}$ 0.063 2.671 15.247 21.587 35.319 0.061
Uniform 0.064 13.619 82.848 50.030 27.068 0.062
$\xi$ $\boldsymbol{\pi}^{J},c_{\nu}=0$ 0.136 0.035 0.017 0.012 0.344 0.017
$\boldsymbol{\pi}^{J},c_{\nu}=1$ 0.148 0.035 0.016 0.011 0.333 0.013
$\boldsymbol{\pi}^{B}$ 0.104 0.028 0.018 0.013 0.373 0.012
$\boldsymbol{\pi}^{D}$ 0.105 0.029 0.019 0.014 0.401 0.013
Uniform 0.105 0.026 0.016 0.010 0.261 0.011
$\omega$ $\boldsymbol{\pi}^{J},c_{\nu}=0$ 0.120 0.071 0.049 0.041 0.156 0.088
$\boldsymbol{\pi}^{J},c_{\nu}=1$ 0.128 0.067 0.046 0.039 0.143 0.085
$\boldsymbol{\pi}^{B}$ 0.099 0.076 0.076 0.064 0.199 0.085
$\boldsymbol{\pi}^{D}$ 0.099 0.078 0.079 0.068 0.211 0.086
Uniform 0.099 0.070 0.047 0.040 0.109 0.083
Table 2: Comparison of the mean absolute deviation (MAD) among $\boldsymbol{\pi}^{J}$ with $c_{\nu}=0$ and $c_{\nu}=1$, $\boldsymbol{\pi}^{B}$, $\boldsymbol{\pi}^{D}$, and the Uniform prior. The maximum a posteriori (MAP) estimation with a Uniform prior distribution is generally equivalent to the maximum likelihood estimation (MLE). Each entry in the table represents the MAD, calculated based on 1000 replications.
3.2 Comparison with Existing Approaches
To evaluate the performance of BIGPAST in comparison to the $z$-score, $t$-score (Crawford and Howell, 1998), Crawford-Garthwaite Bayesian approach (Crawford and Garthwaite, 2007), and Anderson-Darling non-parametric approach (Scholz and Stephens, 1987) for abnormality detection in a single-subject against a control group, we adopt the “severe skew ($\gamma_{1}=-0.7$)” settings from Section 3.1.1 in Crawford et al. (2006). The parameter $\gamma_{1}$ is a measure of skewness in Crawford et al. (2006). This setting is equivalent to parameter settings: $\alpha=-3.23$, $\nu=7$, $\xi=0$, and $\omega=1$ in the definition of our skewed Student’s $t$ distribution. The single-subject and control group observations are generated from $f(x|-3.23,7,0,1)$. We consider sample sizes of 50, 100, 200, and 400 for the control group. The code for the simulations in this section can be found in GitHub repository: BIGPAST. Given a significance level of 0.05 and a one-sided test (less), we record the abnormality detection results using the $z$-score, $t$-score, Crawford-Garthwaite Bayesian method, Anderson-Darling method, and BIGPAST. This procedure is repeated one million times as Crawford et al. (2006). The results, presented in Table 3, indicate that BIGPAST outperforms the other methods in terms of false positive rate (FPR) and accuracy (ACC) when all single-subject observations are from the ground true distribution.
$n$ $z$ $t$ CG AD BIGPAST
FPR 50 0.070022 0.066415 0.066427 0.073271 0.051883
100 0.069632 0.067787 0.067786 0.056572 0.053742
200 0.066616 0.065660 0.065668 0.057564 0.052586
400 0.067888 0.067435 0.067436 0.056244 0.052051
ACC 50 0.929978 0.933585 0.933573 0.926729 0.948117
100 0.930368 0.932213 0.932214 0.943428 0.946258
200 0.933384 0.934340 0.934332 0.942436 0.947414
400 0.932112 0.932565 0.932564 0.943756 0.947949
Table 3: The table presents the simulation results for false positive rate (FPR) and accuracy (ACC), derived from single-subject observations that are exclusively negative, with one-sided test (less). Each entry in the table is computed based on one million test outcomes. The abbreviations ‘CG’ and ‘AD’ denote the Crawford-Garthwaite Bayesian and Anderson-Darling methods, respectively.
Given that all single-subject observations are from ground true distribution, implying a negative underlying result, the true positive rate and false negative rate (Type II error) are undefined. To address this, we modify the generation of single-subject observations such that 50% yield negative results and the remaining 50% yield positive results. These observations are generated using the method described in Section 3.4. The simulation is then conducted as previously described. The true positive rate and false negative rate for the $z$ score, $t$ score, Crawford-Garthwaite Bayesian method, Anderson-Darling method, and BIGPAST, pertaining to a two-sided test, are presented in Table 4. Similar simulations have been conducted for one-sided test (greater) and one-sided test (less), with the results detailed in Appendix, see Tables 7 and 8.
From Table 4, we can conclude that BIGPAST outperforms the $t$-score and Crawford-Garthwaite methods in terms of false positive rate (FPR) and accuracy (ACC) when the single-subject observations are mixed. As normal distribution is a special case of skewed Student’s $t$ distribution ($\alpha=0,\nu=+\infty$), the performance of BIGPAST is still as good as those of $t$-score and Crawford-Garthwaite methods when the underlying distribution is normal distribution, see Table 9 in Appendix. The results are not surprising because the errors in this context originate from two primary sources: model misspecification and inherent randomness. Notably, BIGPAST is capable of effectively mitigating the errors arising from model misspecification; see Figure 3 in the next section.
$n$ $z$ $t$ CG AD BIGPAST
FPR 50 0.1320 0.1182 0.1182 0.1107 0.0428
100 0.1269 0.1199 0.1199 0.0782 0.0349
200 0.1207 0.1166 0.1167 0.0644 0.0334
400 0.1236 0.1215 0.1215 0.0368 0.0187
TPR 50 0.6908 0.6502 0.6502 0.9101 0.8327
100 0.6597 0.6415 0.6416 0.9356 0.8992
200 0.6420 0.6312 0.6311 0.9655 0.9552
400 0.6316 0.6264 0.6264 0.9660 0.9413
ACC 50 0.7794 0.7660 0.7660 0.8997 0.8949
100 0.7664 0.7608 0.7609 0.9287 0.9322
200 0.7606 0.7573 0.7572 0.9505 0.9609
400 0.7540 0.7525 0.7524 0.9646 0.9613
Table 4: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under two-sided test. Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’ and ‘AD’ represent the Crawford-Garthwaite Bayesian and the Anderson-Darling methods, respectively.
3.3 Model Misspecification Error
In this section, we conduct a comparative analysis of the BIGPAST, CG, and AD methodologies under varying alternative hypotheses and parameter configurations. The predetermined significance level is 0.05. The skewness parameter $\alpha$ and degree of freedom $\nu$ are assigned values from the sets $(1,10)$, $(1,5)$, $(2,5)$, and $(3,5)$, respectively. The location and scale of skewed Student’s $t$ distribution are set to 0 and 1, respectively. Other sample settings include a sample size $n=100$, $c_{\nu}=1$, and a Metropolis-Hastings sampling size of 2000. The comparison procedures are as follows: Firstly, $N=200$ independent samples are randomly generated from the skewed Student’s $t$ distribution with parameters $\alpha$ and $\nu$. Secondly, $m=1000$ single-subjects are randomly drawn, and the BIGPAST, CG, and AD tests are conducted on each of the $N$ samples. Thirdly, for each of $N$ samples, the Type I error (false positive rate), power (true positive rate), and accuracy are calculated based on the $m=1000$ single-subjects. Finally, the Type I error, power, and accuracy are reported in the form of boxplots for the four skewed Student’s $t$ parameter settings, as shown in Figures 10, 11, 12 and 2.
Under the mechanism of these parameter settings, the model is misspecified for the CG test. The theoretical assumption of CG test is the normal distribution, the mean and variance can be directly computed from the skewed Student’s $t$ distribution by $\mu=\delta b_{\nu},\nu>1$ and $\sigma^{2}=\nu/(\nu-2)-(b_{\nu}\delta)^{2},\nu>2$, where
$b_{\nu}=\frac{\sqrt{\nu}\Gamma\left(\frac{1}{2}(\nu-1)\right)}{\sqrt{\pi}\Gamma\left(\frac{\nu}{2}\right)},\quad\nu>1;\quad\delta=\frac{\alpha}{\sqrt{1+\alpha^{2}}}.$
We utilise the total variation distance to quantify the disparity between skewed Student’s $t$ (for BIGPAST) and normal (for CG) distributions. The total variation distances of the aforementioned four settings are: 0.057, 0.111, 0.155 and 0.186, respectively, as illustrated in Figures 10(a), 11(a), 12(a) and 2(a). In the context of a two-sided hypothesis test, the divergence between the BIGPAST and CG methodologies becomes larger as the total variation distance increases, as shown in Figures 10, 11, 12 and 2. The results of the AD approach demonstrate robustness for different total variation distances, but perform significantly worse than BIGPAST in terms of Type I error, as seen in Figure 2(b). The AD approach also performs slightly worse than BIGPAST in terms of accuracy, as depicted in Figure 2(d).
(a) The density
(b) The Type I error
(c) The power
(d) The accuracy
Figure 2: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=3$ and $\nu=5$.  Figure 2(a) shows the densities of skewed Student’s $t$ and normal distributions, the $\mu$ and $\sigma^{2}$ of normal distribution are equal to the mean and variance of skewed Student’s $t$ distribution. In fact, the normal distribution is the theoretical assumption of the CG test when the sample comes from the skewed Student’s $t$ distribution. AD test is a nonparametric test; there is one result that is compared in the ‘two-sided’ category. TV distance is short for total variation distance. Each boxplot summarises over 200 independent replications.
The complexity of the one-sided test (either ‘greater’ or ‘less’) compared to the two-sided test due to the skewed Student’s $t$ distribution increases due to skewed Student’s $t$ distribution has light and heavy tails. The model misspecification error impacts the tails of the skewed Student’s $t$ distribution differently. When $\alpha>0$ (resp. $\alpha<0$), the skewed Student’s $t$ distribution has a right (resp. left) heavy tail. In this paragraph, we explore the scenario where the direction of the alternative hypothesis is ‘greater’, i.e., one-sided test (greater). If we implement the CG test, the 95% quantile of normal distribution is less than the 95% quantile of skewed Student’s $t$ distribution. Consequently, the Type I error of the CG test exceeds that of the BIGPAST test, leading to a higher power for the CG test, as shown in Figures 2(b) and 2(c). This can be explained by the light-red region between the 95% quantile of skewed Student’s $t$ distribution and the 95% quantile of normal distribution, as depicted in Figure 3(b). Theoretically, all single-case observations from this region will be classified as positive by the CG test and as negative by the BIGPAST test, leading us to refer to this region as the false positive region. The error generated by the false positive region arises from model misspecification rather than sampling randomness, as illustrated in Figure 3.
Conversely, for the light tail of the skewed Student’s $t$ distribution, the 5% quantile of the normal distribution is less than the 5% quantile of the skewed Student’s $t$ distribution. Given that the alternative is ‘less’ and $\alpha>0$, the region between the 5% quantile of skewed Student’s $t$ distribution and the 5% quantile of the normal distribution is referred to as the false negative region other than the false positive region. Because all the single-case observations from this region will be classified as negative by the CG test, but as positive by the BIGPAST test. Consequently, the theoretical Type I error is zero, excluding the random error of sampling, as depicted in Figure 2(b). The power of the CG test, i.e., the actual positive rate, is lower than that of the BIGPAST test, as illustrated in Figure 2(c).
(a) Alternative: ‘less’
(b) Alternative: ‘greater’
Figure 3: Model misspecification error on the light and heavy tails for one-sided test.
For the left skewed Student’s $t$ distribution ($\alpha<0$), we can obtain similar results just by reflecting the densities in line $x=0$. This will not be detailed here. From the preceding discussion, we can conclude that (1) The difference between the BIGPAST and CG methodologies becomes more significant in terms of Type I error, power, and accuracy as the total variation distance increases; (2) Assuming a skewed Student’s $t$ distribution, the BIGPAST approach outperforms the CG approach in terms of accuracy due to the presence of model misspecification error; (3) Given the asymmetry between the light and heavy tails of the skewed Student’s $t$ distribution, we recommend using a two-sided test instead of a one-sided test. This is because the combinations of a one-sided test (‘less’ or ‘greater’) and the sign of $\alpha$ significantly influence the evaluation of detection.
3.4 Comparison Study with Other Frameworks
This section is dedicated to assessing the performance of the proposed BIGPAST methodology and existing approaches when a control group is present. The first approach is a variant of BIGPAST based on the maximum a posteriori (MAP) estimation. Specifically, procedures 3-5 of the BIGPAST are simplified into the following two steps:
  1. 3
    Obtain the maximum posterior estimators of $\alpha,\nu,\xi,\omega$ by maximising the posterior distribution $\boldsymbol{\pi}(\alpha,\nu,\xi,\omega|\mathbf{x})$, denoted as $\hat{\alpha}_{\mathrm{MAP}},\hat{\nu}_{\mathrm{MAP}},\hat{\xi}_{\mathrm{MAP}},\hat{\omega}_{\mathrm{MAP}}$.
  2. 4
    Given the significance level $\alpha^{*}$, the credible interval of $x^{*}$ given $\mathbf{x}$ is $[x^{\mathrm{MAP}}_{\alpha^{*}/2},x^{\mathrm{MAP}}_{1-\alpha^{*}/2}]$, where $x^{\mathrm{MAP}}_{\alpha^{*}/2}$ and $x^{\mathrm{MAP}}_{1-\alpha^{*}/2}$ are the $\alpha^{*}/2$ and $21-\alpha^{*}/$ 2 quantiles of $\boldsymbol{\pi}(\hat{\alpha}_{\mathrm{MAP}},\hat{\nu}_{\mathrm{MAP}},\hat{\xi}_{\mathrm{MAP}},\hat{\omega}_{\mathrm{MAP}}|\mathbf{x})$, respectively. If $x^{*}$ does not fall within this credible interval, then the null hypothesis $\mathcal{H}_{0}$ can be rejected at the significance level $\alpha^{*}$.
The second approach is a non-Bayesian approach that relies on maximum likelihood estimation (MLE). This approach operates under the sole assumption that the control group follows a skewed Student’s $t$ distribution. The MLE procedures can be outlined as follows:
  1. 1
    Obtain the maximum likelihood estimators of $\alpha,\nu,\xi,\omega$ by maximising the likelihood function $l(\mathbf{x}|\alpha,\nu,\xi,\omega)$. These estimators are denoted as $\hat{\alpha}_{\mathrm{MLE}},\hat{\nu}_{\mathrm{MLE}},\hat{\xi}_{\mathrm{MLE}},\hat{\omega}_{\mathrm{MLE}}$.
  2. 2
    Given the significance level $\alpha^{*}$, the confidence interval of $x^{*}$ given $\mathbf{x}$ is $[x^{\mathrm{MLE}}_{\alpha^{*}/2},x^{\mathrm{MLE}}_{1-\alpha^{*}/2}]$, where $x^{\mathrm{MLE}}_{\alpha^{*}/2}$ and $x^{\mathrm{MLE}}_{1-\alpha^{*}/2}$ are the $\alpha^{*}/2$ and $21-\alpha^{*}/$ 2 quantiles of $f_{c}(\hat{\alpha}_{\mathrm{MLE}},\hat{\nu}_{\mathrm{MLE}},\hat{\xi}_{\mathrm{MLE}},\hat{\omega}_{\mathrm{MLE}})$, respectively. If $x^{*}$ does not fall within this confidence interval, then the null hypothesis $\mathcal{H}_{0}$ can be rejected at the significance level $\alpha^{*}$.
The third approach is a non-parametric (NP) approach. Given group data $\mathbf{x}$ and a significance level $\alpha^{*}$, we can derive the empirical confidence interval $[x^{\mathrm{NP}}_{(\lfloor n*\alpha^{*}/2\rfloor)},x^{\mathrm{NP}}_{(\lceil n*(1-\alpha^{*}/2)\rceil)}]$ where $x^{\mathrm{NP}}_{(\lfloor n*\alpha^{*}/2\rfloor)}$ and $x^{\mathrm{NP}}_{(\lceil n*(1-\alpha^{*}/2)\rceil)}$ are the empirical quantiles of $\mathbf{x}$ at $\alpha^{*}/2$ and $21-\alpha^{*}/$ 2, respectively. If $x^{*}$ does not fall within this empirical confidence interval, then the null hypothesis $\mathcal{H}_{0}$ can be rejected at the significance level $\alpha^{*}$. The NP approach does not rely on any distributional assumptions. Given that our BIGPAST methodology employs the Metropolis-Hastings sampler in Step 3, we will refer to the aforementioned four approaches: MH, MAP, MLE, and NP, respectively.
We now outline the data generation mechanism for a single-subject and control group. Given the parameters $\alpha,\nu,\xi,\omega$, we randomly draw $N$ independent control groups $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}$ from $f_{c}(\alpha,\nu,\xi,\omega)$. The following is a straightforward method for generating single-subject data. Define $S_{1}\coloneqq(-\infty,q_{2.5\%}]\cup[q_{97.5\%},\infty)$ and $S_{2}\coloneqq(q_{2.5\%},q_{5\%}]\cup[q_{95\%},q_{97.5\%})$ where $q_{2.5\%},q_{5\%},q_{95\%},q_{97.5\%}$ are 0.025, 0.05, 0.95, 0.975 quantiles of $f_{c}(\alpha,\nu,\xi,\omega)$. We draw $k_{1}$ and $k_{2}$ random copies of single-subjects $x^{*}_{1}\ldots,x^{*}_{k_{1}},x^{*}_{k_{1}+1},\ldots,x^{*}_{k_{1}+k_{2}}$ respectively from the uniform distributions defined on $S_{1}$ and $S_{2}$. Let $K=k_{1}+k_{2}$. At the significance level of 0.05, the ground truth for the first $k_{1}$ single-subjects is negative, while the ground truth for the remaining $k_{2}$ single-subjects is positive. In the simulation, we first apply the MH, MAP, MLE, and NP approaches to $\mathbf{x}_{i},i=1,2,\ldots,N$ to obtain the four types of intervals. Next, we conduct the hypothesis tests by comparing $\mathbf{x}^{*}=(x^{*}_{1},x^{*}_{2},\ldots,x^{*}_{K})$ with each type of intervals. Finally, we compute the false discovery rate (FDR) and accuracy (ACC) defined by
$\mathrm{FDR}_{i}^{\mathcal{M}}\coloneqq\frac{\mathrm{FP}_{i}^{\mathcal{M}}}{\mathrm{TP}_{i}^{\mathcal{M}}+\mathrm{FP}_{i}^{\mathcal{M}}},\quad\mathrm{ACC}_{i}^{\mathcal{M}}\coloneqq\frac{\mathrm{TP}_{i}^{\mathcal{M}}+\mathrm{TN}_{i}^{\mathcal{M}}}{K},\quad i=1,2,\ldots,N,$
where $\mathrm{TP}$, $\mathrm{FP}$, and $\mathrm{TN}$ represent the number of true positive, false positive and true negative, respectively. $\mathcal{M}$ could be one of MH, MAP, MLE, NP.
The parameters settings are as follows: we consider ten combinations of $(\alpha,\nu)$: (0,3), (1,3), (3,3), (5,5), (10,10), (20,20), (30,30), (50,50), (5,50) and (50,5) with fixed $\xi=-2$ and $\omega=2$. In MH approach, we let $m=5000$, $s=100$ and $c_{\nu}=1$. The other parameter settings are $N=400$, $n=100$, $k_{1}=k_{2}=1000$, $\alpha^{*}=0.05$. The results are presented in Table 5 and Figure 4.
Table 5 displays the FDR and ACC of the four approaches. The MH and MAP approaches consistently outperform the NP and MLE approaches in terms of FDR and ACC. When ${\left\lvert\alpha\right\rvert}\leq 3$ and $\nu\leq 3$, both the FDR and ACC of the MAP approach surpass those of the MH approach. This is expected, as the MAP approach provides more accurate estimation of the underlying parameters of the skewed Student’s $t$ distribution when ${\left\lvert\alpha\right\rvert}$ and $\nu$ are small, as is shown in column 3 of Table 2. When ${\left\lvert\alpha\right\rvert}$ or $\nu$ is large, the FDR and ACC of the MH approach perform better than those of the MAP approach. This can be attributed to the randomness of parameters in the Metropolis-Hastings sampler. However, the MAP approach fails to estimate the underlying parameters when $\alpha$ or $\nu$ is large, as is shown in column 4 of Table 2, and lacks randomness of parameters. The boxplots in Figure 4 clearly illustrate the comparison between the MH and MAP approaches. The results illustrate that the FDR of MH (i.e. BIGPAST) is generally lower than others, and the ACC of MH (i.e. BIGPAST) is generally greater than other methods in conducting hypothesis tests for single-subject data.
$(\alpha,\nu)$ FDR ACC
MH MAP NP MLE MH MAP NP MLE
(0,3) 0.174 0.165 0.216 0.266 0.865 0.872 0.833 0.787
(1,3) 0.158 0.145 0.201 0.257 0.877 0.888 0.845 0.797
(3,3) 0.142 0.129 0.197 0.254 0.891 0.901 0.849 0.798
(5,5) 0.131 0.136 0.206 0.255 0.905 0.899 0.846 0.803
(10,10) 0.118 0.145 0.200 0.259 0.915 0.894 0.851 0.797
(20,20) 0.132 0.157 0.190 0.255 0.907 0.886 0.861 0.803
(30,30) 0.132 0.171 0.188 0.261 0.907 0.875 0.862 0.796
(50,50) 0.134 0.184 0.193 0.254 0.906 0.864 0.859 0.801
(5,50) 0.144 0.158 0.185 0.261 0.898 0.885 0.865 0.796
(50,5) 0.139 0.165 0.218 0.249 0.899 0.876 0.834 0.805
Table 5: The average false discovery rate (FDR) and accuracy (ACC) for the MH, MAP, MLE, and NP approaches, calculated over 400 repetitions. In this context, MH refers to the original BIGPAST method, MAP denotes the Bayesian inference framework based on maximum a posteriori estimation, MLE represents the method based on maximum likelihood estimation, and NP signifies the non-parametric method.
(a) $\alpha=0,\nu=3$
(b) $\alpha=5,\nu=5$
(c) $\alpha=50,\nu=50$
Figure 4: The boxplots for false discovery rate (FDR) and accuracy (ACC) of the MH, MAP, MLE, and NP approaches for three distinct skewness parameters: low ($\alpha=0$), medium ($\alpha=5$), and high ($\alpha=50$).In this context, MH refers to the original BIGPAST method, MAP denotes the Bayesian inference framework based on maximum a posteriori estimation, MLE represents the method based on maximum likelihood estimation, and NP signifies the non-parametric method.
4 Real Data Analysis
This dataset is derived from a mild traumatic brain injury (mTBI) study conducted by Innovision IP Ltd. It comprises functional time series of Magnetoencephalography (MEG) data and an anatomical T1-weighted Magnetic Resonance Imaging (MRI) scan for 103 healthy subjects and one patient with mild traumatic brain injury. The MEG data, recorded from 306 sensors, was pre-processed using the ‘MNE-Python’ package (Gramfort et al., 2013). The MEG data was analysed to reconstruct the source generators using a minimum norm inverse method. The MRI data was pre-processed with the ‘FreeSurfer’ package (Han et al., 2006). The standard FreeSurfer pipeline (detailed on their website) was followed and includes motion correction, intensity normalisation, skull-stripping, normalisation to template space and segmentation of the cortex and subcortical structures using the Desikan-Killiany atlas. This results in 68 regions covering the entire cortex. For each cortical region, we calculate the average of source powers within this region. This process yields a control group dataset with dimensions $103\times 68\times 100$ and a patient dataset with dimensions $68\times 100$. To simplify, we treat the epochs as repeated observations and take average over them to reduce the data’s complexity. Consequently, the control group dataset has dimensions $103\times 68$, and the patient dataset is reduced to a vector of length $68$.
For each brain region of the Desikan-Killiany cortical atlas, we apply the BIGPAST method to the control group and patient dataset to test whether the patient significantly differs from the control group. In this real data analysis, we know that the single-subject has mTBI. However, in practice, we can calculate the observation of a new client and conduct the hypothesis test using BIGPAST to determine whether the client significantly differs from the control group.
We further explore the impact of waveband frequency on the comparison between a control group and a single-subject. The wavebands are arranged in ascending order of frequency: delta (1-4 Hz), theta (4-8 Hz), alpha (8-12 Hz), beta (12-30 Hz), and gamma (30-45 Hz). The subsequent results also include their findings, denoted as CG for brevity. The prespecified significance level is 0.05.
(CG, BIGPAST) delta theta alpha beta gamma
(mTBI, mTBI)
6,17,
40,51
4,6,
17,48
6,14,
17,40
6,12,17,
38,40
17
(Normal, mTBI)
1,10,16,26,
35,50,59,62
1,2,24,
35,62
2,24,35,
58,59
1,2,10,22,
24,35,59
1,2,35
(mTBI, Normal)
NA
14,40 NA
14,47
6,40
(Normal, Normal) Others Others Others Others Others
Table 6: The test results of the Crawford-Garthwaite (CG) and BIGPAST approaches, both based on the normal and skewed Student’s $t$ assumptions. The indices corresponding to specific brain cortical regions are 1: lh-caudal anterior cingulate, 2: lh-caudal middle frontal, 4: lh-entorhinal, 6: lh-fusiform, 10: lh-isthums cingulate, 12: lh-lateral orbitofrontal, 14: lh-medial orbitofrontal, 17: lh-parahippocampal, 22: lh-postcentral, 24: lh-precentral, 26: lh-rostral anterior cingulate, 35: rh-caudal anterior cingulate, 38: rh-entorhinal, 40: rh-fusiform, 47: rh-lingual, 50: rh-paracentral, 51: rh-parahippocampal, 58: rh-precentral, 59: rh-precuneus, 62: rh-superior frontal. The prefixes ‘lh-’ and ‘rh-’ indicate the left and right hemispheres of the brain, respectively. ‘NA’ stands for ‘not available’, and ‘Others’ represents the remaining brain cortical indices not listed in the column, but can be found the indices in Table 10.
(a) Delta Band, $\bar{\hat{\alpha}}\approx 3.95$, Left Caudal Anterior Cingulate
(b) Gamma Band, $\bar{\hat{\alpha}}\approx 2.69$, Right Fusiform
Figure 5: The comparison results between the Crawford-Garthwaite (CG) and BIGPAST approaches in terms of density estimations. The light-red lines represent the skewed Student’s $t$ densities, with parameters drawn from the BIGPAST approach’s posterior distribution. The light green lines depict the normal densities, with parameters drawn from the CG approach’s posterior distribution. The CG test is based on 10,000 densities, while the BIGPAST test is based on 5,000 densities (the first 5,000 densities are discarded for burn-in). For clarity, each panel only displays 200 densities drawn from the CG and BIGPAST posterior distribution. The red vertical line on the $x$ axis represents the observation of a single-subject (i.e., a patient with mTBI). A histogram is used to demonstrate the goodness of fit for both the CG and BIGPAST approaches.
In this real data analysis, the null hypothesis is two-sided, with one heavy tail and one light tail in the skewed Student’s $t$ distribution. The cortical region indices in row 3 of Table 6 exhibit a positive skewness parameter, i.e., $\alpha>0$. Consequently, the light tail is on the left, and the heavy tail is on the right. For all cortical region indices in row 3 of Table 6, the single-subject observation is located at the light tail of the skewed Student’s $t$ distribution. This positioning suggests a high probability of rejecting the null hypothesis based on the skewed Student’s $t$ assumption, as illustrated in Figure 5(a). Similarly, for cortical regions 6 (lh-fusion, gamma band), 14 (lh-medial orbitofrontal,theta and beta bands), 40 (rh-fusiform, theta and gamma bands), and 47 (rh-lingual, beta band), the single-subject observation is located at the heavy tail of the skewed Student’s $t$ distribution, as shown in Figure 5(b). The skewed Student’s $t$ assumption is more appropriate for these cortical regions than the assumption of normal distribution. The BIGPAST approach outperforms the CG approach in detecting the single-subject observation at the heavy tail of the skewed Student’s $t$ distribution, as demonstrated in Figure 5. Because the BIGPAST approach corrects the model misspecification error at both tails compared to the CG approach. By combining rows 2 and 3 of Table 6 and Figure 5, we conclude that the BIGPAST approach is more reliable than the CG approach in hypothesis tests for single-subject data when the control group data exhibits skewness.
5 Conclusion
In this paper, we proposed a novel Bayesian inference framework for the abnormality detection on a single-subject versus a control group. Under the assumption of a normal distribution, the $t$-score (Crawford and Howell, 1998) is capable of managing scenarios where the sample size is fewer than 30 (Machin et al., 2018). However, in neuroscience, the number of epochs often exceeds 100, sometimes even reaching 1000, and the distributions—whether over epochs or subjects—exhibit skewness. Under these conditions, methods predicated on the assumption of normality cease to be effective. Our proposed methodology, BIGPAST, leverages the skewed Student’s $t$ distribution, effectively capturing the inherent asymmetry of the data.
In terms of accuracy, power, and Type I error, our proposed BIGPAST methodology outperforms the existing Crawford-Garthwaite (CG) approach. Particularly when the control group data exhibits skewness, BIGPAST provides a more reliable solution for hypothesis testing on single-subject data. Furthermore, BIGPAST demonstrates superior robustness to model misspecification errors compared to the CG approach. The performance of the BIGPAST and CG methodologies is significantly influenced by the total variation distance between the skewed Student’s $t$ distribution and the normal distribution. When this distance is large, the BIGPAST approach exhibits superior accuracy compared to the CG method. The skewed Student’s $t$ distribution’s asymmetry, characterised by its light and heavy tails, further emphasises the necessity of employing a two-sided test over a one-sided test.
The BIGPAST framework can be readily adapted to accommodate other distributions, such as Sub-normal, Sub-Weibull, or Extreme Value distributions. The key requirement is to identify an appropriate prior distribution for the parameters of the chosen distribution. A further extension of BIGPAST could encompass multivariate distributions, thereby enabling the application of the BIGPAST methodology to high-dimensional data in future.
Data and Code Availability
The calculation of skewed Student’s $t$ distribution is implemented in the Python package skewt-scipy (https://pypi.org/project/skewt-scipy/). The code for simulation studies can be found in GitHub repository: BIGPAST. The real data comes from the Cambridge Centre for Ageing and Neuroscience (Cam-CAN) dataset (Shafto et al., 2014).
Competing interests
The authors from the University of Kent, J.L., J.Z., and P.L. have no competing interests. G.G. and S.C. are employees of Innovision IP Ltd which provides commercial reports on individuals who may have had a head injury.
Author contributions statement
JL: Conceptualisation, Methodology, Programming and Software, Investigation, and Writing—Original Draft. GG: Conceptualisation, Supervision, Methodology, Resources, Investigation, Project Administration and Writing—Review & Editing. SC: Methodology and Writing- Review & Editing. PL: Supervision, Methodology, Project Administration and Writing—Review & Editing. JZ: Conceptualisation, Supervision, Methodology, Investigation, Project Administration and Writing—Review & Editing.
Acknowledgments
This work was supported by the High Performance Computing Cluster at University of Kent.
Appendix A Appendix
Lemma A.1.
Let $t(z|\nu)$ and $T(z|\nu)$ be the probability density function and cumulative distribution function of the Student’s $t$-distribution, respectively, for $\nu\geq 1$, then $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ can be approximated by $t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$, where $\sigma_{\nu}=1.5536$ if $\nu>2400$. If $1\leq\nu\leq 2400$, then
$\sigma_{\nu}$ $=0.00000543(\log(\nu))^{7}-0.00016303(\log(\nu))^{6}+0.00199613(\log(\nu))^{5}-0.01285016(\log(\nu))^{4}$ (5)
$\quad+0.04631303(\log(\nu))^{3}-0.08761023(\log(\nu))^{2}+0.05036188\log(\nu)+1.62021189$
Proof.
Firstly, when $1\leq\nu\leq 2400$, we can use $t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$ to approximate $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ by the similar discussion in Bayes and Branco (2007). Apparently, the optimal scale $\sigma_{\nu}$ depends on $\nu$. To obtain Equation 5, for simplicity, let $f\coloneqq t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ and $g\coloneqq t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$. Define the total variational distance between $f$ and $g$ as
$TV(\sigma_{\nu})\coloneqq TV(f,g)=\frac{1}{2}\int_{-\infty}^{\infty}{\left\lvert f-g\right\rvert}\,\mathrm{d}z$ (6)
Let
$\nu_{i}=\begin{cases}0.5+0.5i&i=1,2,\ldots,199\\ 100+50*(i-199)&i=200,201,\ldots,245,\end{cases}$
given $\nu_{i}$, we use grid search method to minimise $KL(\sigma(\nu_{i}))$ and obtain the optimal scale $\hat{\sigma}(\nu_{i})$. Let $\sigma_{j,\nu_{i}}=1+0.0002j$, $j=0,1,\ldots,5000$, then $\hat{\sigma}_{\nu_{i}}=\operatorname*{arg\,min}_{\sigma_{j,\nu_{i}}}KL(\sigma_{j,\nu_{i}})$ where $KL(\sigma_{j,\nu_{i}})$ is evaluated using the Monte Carlo method. Finally, we carry out the polynomial fit with polynomial degree $p=2,3,\ldots,7$ based on the optimal scale $\hat{\sigma}(\nu_{i})$ and $\log(\nu_{i})$, $i=1,2,\ldots,245$Equation 5 is the best polynomial fit with polynomial degree $p=7$ based on Bayes information criterion, see Figure 6. Secondly, when $\nu>2400$, $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ can be approximated by $\phi(z)[\Phi(z)]^{-1/2}[1-\Phi(z)]^{-1/2}/\pi$. According to Bayes and Branco (2007), $\phi(z)[\Phi(z)]^{-1/2}[1-\Phi(z)]^{-1/2}/\pi$ can be further approximated by $N(0,\sigma^{2}_{\nu})$ where the optimal scale $\sigma_{\nu}$ chosen by total variational distance is 1.5536. Note that Bayes and Branco (2007) used the optimal scale $\sigma_{\nu}$ chosen by the Kullback-Leibler divergence (defined in Equation 7), the optimal scale $\sigma_{\nu}$ is 1.54 for normal distribution. However, our simulation result shows that the total variational distance criterion performs better than the Kullback-Leibler divergence criterion for Student’s $t$ with a small degree of freedom, see Figure 7. Therefore, we employ the criterion of total variational distance to choose the optimal scale $\sigma_{\nu}$ in $t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$ to approximate $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ throughout this paper. Figure 8 shows the density $t(z|\nu)[T(z|\nu)]^{-1/2}[1-T(z|\nu)]^{-1/2}/\pi$ and its approximation $t(z/\sigma_{\nu}|\nu)/\sigma_{\nu}$ with $\sigma_{\nu}$ in Equation 5 for $\nu=1,5,10,100$.
Figure 6: Polynomial fitted curve
Figure 7: Comparison of approximation based on Kullback-Leibler divergence and total variational distance when the degree of freedom $\nu=2$. The blue line represents the density of $f$, the green line represents the density of approximation $g$ based on total variational distance, the orange line is the density of approximation $g$ based on Kullback-Leibler divergence.
[…]
(a) $\nu=1$
(b) $\nu=5$
(c) $\nu=10$
(d) $\nu=100$
Figure 8: Density $f$ (solid blue curve) and its approximation $g$ with $\sigma_{\nu}$ (dashed orange curve) for $\nu=1,5,10,100$.
Lemma A.2.
Let $f(z|\alpha,\nu)$ be the non-linear skewed Student’s $t$ density function, defined as follows:
$f(z|\alpha,\nu)\coloneqq 2t(z|\nu)T(\alpha zr(z,\nu)|\nu+1),\quad z\in\mathbb{R},\alpha\in\mathbb{R},\nu>0,$
where $r(z,\nu)=\sqrt{(\nu+1)/(\nu+z^{2})}$. Given any $\nu>0$, let $\boldsymbol{\pi}(\alpha|\nu)$ represent the Jeffery’s prior for $\alpha$, we have
$\boldsymbol{\pi}(\alpha|\nu)\propto\frac{\sqrt{\pi}\Gamma\left(\frac{v}{2}+1\right)\sqrt{{}_{2}F_{1}\left(\frac{1}{2},v+1;\frac{v+1}{2};-\frac{\alpha^{2}}{\sigma^{2}_{\nu+1}}\right)-\,_{2}F_{1}\left(\frac{1}{2},v+2;\frac{v+1}{2};-\frac{\alpha^{2}}{\sigma^{2}_{\nu+1}}\right)}}{{\left\lvert\alpha\right\rvert}\Gamma\left(\frac{v+1}{2}\right)},$ (8)
where ${}_{2}F_{1}\left(a,b;c;d\right)$ represents the hypergeometric function, $\sigma_{\nu+1}$ is defined in Lemma A.1, $\Gamma(\cdot)$ is the Gamma function.
Proof.
By definition of Jeffery’s prior, we have $\boldsymbol{\pi}(\alpha|\nu)\propto\sqrt{I_{\alpha\alpha}}$, where $I_{\alpha\alpha}$ is the Fisher information of $\alpha$ in the non-linear skewed Student’s $t$ density function $f(z|\alpha,\nu)$. According to Branco et al. (2013, Proposition 2), the Fisher information of $\alpha$ in $f(z|\alpha,\nu)$ is given by
$I_{\alpha\alpha}=2\int_{0}^{+\infty}\frac{z^{2}r(z,\nu)^{2}t(z|\nu)t^{2}(\alpha zr(z,\nu)|\nu+1)}{T(\alpha zr(z,\nu)|\nu+1)\left[1-T(\alpha zr(z,\nu)|\nu+1)\right]}\,\mathrm{d}z.$ (9)
Let $m=\alpha zr(z,\nu)$, by the change of variable $z=\frac{\sqrt{m^{2}\nu}}{\sqrt{\alpha^{2}\nu+\alpha^{2}-m^{2}}}$quation 9 can be rewritten as
$I_{\alpha\alpha}=2\int_{0}^{\alpha\sqrt{\nu+1}}\frac{m^{2}t\left(\frac{\sqrt{m^{2}\nu}}{\sqrt{\alpha^{2}\nu+\alpha^{2}-m^{2}}}\Big{|}\nu\right)t^{2}(m|\nu+1)}{T(m|\nu+1)\left[1-T(m|\nu+1)\right]}\frac{\sqrt{\nu}(\nu+1)}{\left(\alpha^{2}(\nu+1)-m^{2}\right)^{3/2}}\,\mathrm{d}m.$ (10)
By Lemma A.1, we have $t(m|\nu+1)[T(m|\nu+1)]^{-1/2}[1-T(m|\nu+1)]^{-1/2}/\pi\approx t(m/\sigma_{\nu+1}|\nu+1)/\sigma_{\nu+1}$. Substituting this approximation into Equation 10, we obtain
$I_{\alpha\alpha}\approx 2\int_{0}^{\alpha\sqrt{\nu+1}}\frac{m^{2}\pi^{2}\left(\frac{1}{\frac{m^{2}}{\alpha^{2}\nu+\alpha^{2}-m^{2}}+1}\right)^{\frac{\nu+1}{2}}\left(\frac{\nu+1}{\frac{m^{2}}{\sigma^{2}_{\nu+1}}+\nu+1}\right)^{\nu+2}}{\left(\alpha^{2}(\nu+1)-m^{2}\right)^{3/2}\sigma^{2}_{\nu+1}B\left(\frac{\nu}{2},\frac{1}{2}\right)B\left(\frac{\nu+1}{2},\frac{1}{2}\right)^{2}}\,\mathrm{d}m,$ (11)
where $B(\cdot,\cdot)$ is the Beta function. Furthermore, let $x=m^{2}$, by change of variable, Equation 11 can be rewritten as
$I_{\alpha\alpha}\approx 2\int_{0}^{\alpha^{2}(\nu+1)}\frac{\pi^{2}\sqrt{x}\left(\frac{1}{\frac{x}{\alpha^{2}(\nu+1)-x}+1}\right)^{\frac{\nu+1}{2}}\left(\frac{\nu+1}{\nu+\frac{x}{\sigma^{2}_{\nu+1}}+1}\right)^{\nu+2}}{\sigma^{2}_{\nu+1}B\left(\frac{\nu}{2},\frac{1}{2}\right)B\left(\frac{\nu+1}{2},\frac{1}{2}\right)^{2}\left(\alpha^{2}(v+1)-x\right)^{3/2}}\,\mathrm{d}x.$ (12)
By integration, we obtain the following expression:
$I_{\alpha\alpha}\approx\frac{\pi\Gamma\left(\frac{\nu}{2}+1\right)^{2}\left(\,{}_{2}F_{1}\left(\frac{1}{2},\nu+1;\frac{\nu+1}{2};-\frac{\alpha^{2}}{\sigma^{2}_{\nu+1}}\right)-\,_{2}F_{1}\left(\frac{1}{2},\nu+2;\frac{\nu+1}{2};-\frac{\alpha^{2}}{\sigma^{2}_{\nu+1}}\right)\right)}{\alpha^{2}\Gamma\left(\frac{\nu+1}{2}\right)^{2}},$
which completes the proof. ∎
The Proof of Theorem 2.1
Proof.
The log-likelihood function of $f(z|\alpha,\nu)$, say $\ell(\alpha,\nu)$, can be expressed as
$\ell(\alpha,\nu)=\log\left(\frac{2}{\sqrt{\pi}}\right)+\frac{\nu+1}{2}\log\left(\frac{\nu}{\nu+z^{2}}\right)-\frac{1}{2}\log(\nu)+\log\left(\Gamma\left(\frac{\nu+1}{2}\right)\right)-\log\left(\Gamma\left(\frac{\nu}{2}\right)\right)+\log\left(T(w|\nu+1)\right).$ (13)
The derivative of $\ell(\alpha,\nu)$ with respect to $\nu$ is given by
$\frac{\partial\ell(\alpha,\nu)}{\partial\nu}=\frac{1}{2}\log\left(\frac{\nu}{\nu+z^{2}}\right)-\frac{1}{2}\psi\left(\frac{\nu}{2}\right)+\frac{1}{2}\psi\left(\frac{\nu+1}{2}\right)+\frac{z^{2}-1}{2\left(\nu+z^{2}\right)}+\frac{1}{T(w|\nu+1)}\frac{\partial T(w|\nu+1)}{\partial\nu},$ (14)
where $\psi(x)\coloneqq\mathrm{d}\log\Gamma(x)/\mathrm{d}x$ is the digamma function. The derivative of $T(w|\nu+1)$ with respect to the parameter $\nu$ can be expressed by
$\frac{\partial T(w|\nu+1)}{\partial\nu}$ $=t(w|\nu+1)\frac{\partial w}{\partial\nu}+\int_{-\infty}^{w}\frac{\partial t(x|\nu+1)}{\partial\nu}\,\mathrm{d}x$ (15)
$=t(w|\nu+1)\frac{\alpha z\left(z^{2}-1\right)}{2\sqrt{\nu+1}\left(\nu+z^{2}\right)^{3/2}}+\frac{1}{2}\left[-\psi\left(\frac{\nu+1}{2}\right)+\psi\left(\frac{\nu}{2}+1\right)\right]T(w|\nu+1)$
$\quad+\int_{-\infty}^{w}\frac{1}{2}\left[\log\left(\frac{\nu+1}{\nu+x^{2}+1}\right)+\frac{x^{2}-1}{\left(\nu+x^{2}+1\right)}\right]t(x|\nu+1)\,\mathrm{d}x$
$=t(w|\nu+1)\frac{\alpha z\left(z^{2}-1\right)}{2\sqrt{\nu+1}\left(\nu+z^{2}\right)^{3/2}}+\frac{1}{2}\left[-\psi\left(\frac{\nu+1}{2}\right)+\psi\left(\frac{\nu}{2}+1\right)\right]T(w|\nu+1)$
$\quad+g(\nu,w)-\frac{w}{2(\nu+1)}t(w|\nu+1)$
$=-t(w|\nu+1)\frac{\alpha z\sqrt{\nu+1}}{2\left(\nu+z^{2}\right)^{3/2}}+\frac{1}{2}\left[-\psi\left(\frac{\nu+1}{2}\right)+\psi\left(\frac{\nu}{2}+1\right)\right]T(w|\nu+1)$
$\quad+g(\nu,w)$
where $g(\nu,w)\coloneqq\int_{-\infty}^{w}\frac{1}{2}\log\left(\frac{\nu+1}{\nu+x^{2}+1}\right)t(x|\nu+1)\,\mathrm{d}x$. Substituting Equation 15 into Equation 14, we obtain
$\frac{\partial\ell(\alpha,\nu)}{\partial\nu}$ $=\frac{1}{2}\log\left(\frac{\nu}{\nu+z^{2}}\right)-\frac{1}{2}\psi\left(\frac{\nu}{2}\right)+\frac{1}{2}\psi\left(\frac{\nu+2}{2}\right)+\frac{z^{2}-1}{2\left(\nu+z^{2}\right)}$ (16)
$\quad-h(w)\frac{\alpha z\sqrt{\nu+1}}{2\left(\nu+z^{2}\right)^{3/2}}+\frac{g(\nu,w)}{T(w|\nu+1)},$
where $h(w)\coloneqq t(w|\nu+1)/T(w|\nu+1)$. Next, we focus on the calculation of $g(\nu,w)$. Note that $\log\frac{\nu+1}{\nu+x^{2}+1}$ and $t(x|\nu+1)$ are symmetric with respect to $x$, therefore for $w>0$, we only consider
$g_{0}(\nu,w)$ $=\int_{0}^{w}\frac{1}{2}\log\left(\frac{\nu+1}{\nu+x^{2}+1}\right)t(x|\nu+1)\,\mathrm{d}x$ (17)
$=\frac{1}{2}\int_{0}^{w}\log\left(\frac{\nu+1}{\nu+x^{2}+1}\right)\frac{\left(\frac{\nu+1}{\nu+x^{2}+1}\right)^{\frac{\nu+2}{2}}}{\sqrt{\nu+1}B\left(\frac{\nu+1}{2},\frac{1}{2}\right)}\,\mathrm{d}x$
$=\frac{1}{4}\int_{0}^{w^{2}}\frac{1}{\sqrt{u}}\log\left(\frac{\nu+1}{\nu+u+1}\right)\frac{\left(\frac{\nu+1}{\nu+u+1}\right)^{\frac{\nu+2}{2}}}{\sqrt{\nu+1}B\left(\frac{\nu+1}{2},\frac{1}{2}\right)}\,\mathrm{d}u$
$=\frac{1}{4B\left(\frac{\nu+1}{2},\frac{1}{2}\right)}\int_{\frac{\nu+1}{\nu+w^{2}+1}}^{1}\frac{y^{\frac{\nu-1}{2}}}{\sqrt{1-y}}\log\left(y\right)\,\mathrm{d}y$
$=+\frac{\sqrt{1+w^{2}+\nu}}{(1+\nu)^{2}}t(w|\nu+1)\,_{3}F_{2}\left(\frac{1}{2},\frac{\nu}{2}+\frac{1}{2},\frac{\nu}{2}+\frac{1}{2};\frac{\nu}{2}+\frac{3}{2},\frac{\nu}{2}+\frac{3}{2};\frac{\nu+1}{\nu+w^{2}+1}\right)$
$\quad-\frac{\sqrt{1+w^{2}+\nu}}{2(1+\nu)}\log\left(\frac{\nu+1}{\nu+w^{2}+1}\right)t(w|\nu+1)\,_{2}F_{1}\left(\frac{1}{2},\frac{\nu+1}{2};\frac{\nu+3}{2};\frac{\nu+1}{\nu+w^{2}+1}\right)$
$\quad-\frac{1}{4}\left[\psi\left(\frac{\nu}{2}+1\right)-\psi\left(\frac{\nu+1}{2}\right)\right].$
Similarly, we have
$g_{1}(\nu)=\int_{0}^{+\infty}\frac{1}{2}\log\left(\frac{\nu+1}{\nu+x^{2}+1}\right)t(x|\nu+1)\,\mathrm{d}x=-\frac{1}{4}\left[\psi\left(\frac{\nu}{2}+1\right)-\psi\left(\frac{\nu+1}{2}\right)\right].$ (18)
Finally, we obtain $g(\nu,w)$ as follows:
$g(\nu,w)=\begin{cases}g_{0}(\nu,w)+g_{1}(\nu)&\text{if }w\geq 0\\ g_{1}(\nu)-g_{0}(\nu,w)&\text{otherwise }.\end{cases}$ (19)
By take the partial derivative of Equation 16 with respect to $\alpha$, we have
$\frac{\partial^{2}\ell(\alpha,\nu)}{\partial\nu\partial\alpha}=-\frac{\alpha z^{2}(\nu+1)}{2(\nu+z^{2})^{2}}h^{\prime}(w)-\frac{z\sqrt{\nu+1}}{2\left(\nu+z^{2}\right)^{3/2}}h(w)+zh(w)\sqrt{\frac{\nu+1}{\nu+z^{2}}}\left[\frac{1}{2}\log\left(\frac{\nu+1}{\nu+w^{2}+1}\right)-\frac{g(\nu,w)}{T(w|\nu+1)}\right],$ (20)
where
$h^{\prime}(w)=-\frac{w(2+\nu)}{1+w^{2}+\nu}h(w)-h^{2}(w).$
Because Equation 20 includes the terms with $z^{1}$ which is odd, therefore the Fisher Information of Equation 20 can be expressed as
$I_{\alpha\nu}=-\mathbb{E}\left[h^{2}(w)\frac{\alpha z^{2}(\nu+1)}{2(\nu+z^{2})^{2}}\right]+\mathbb{E}\left[zh(w)\sqrt{\frac{\nu+1}{\nu+z^{2}}}\frac{g(\nu,w)}{T(w|\nu+1)}\right]$ (21)
By numeric results, we found that
$\frac{g(\nu,w)}{T(w|\nu+1)}=2c_{\nu}*g_{1}(\nu),$
where $c_{\nu}$ is a constant only depending on $\nu$, see Figure 9.
(a) $\nu=1$
(b) $\nu=5$
(c) $\nu=10$
(d) $\nu=100$
Figure 9: The $c_{\nu}$ for $\nu=1,5,10,100$. When $w\rightarrow-\infty$, $c_{\nu}\rightarrow 0$; When $w\rightarrow\infty$, $c_{\nu}\rightarrow 1$. Specially, if $w=0$, $c_{\nu}=1/2$. In the above figures, $c_{\nu}<160$.
Therefore, Equation 21 can be written as
$I_{\alpha\nu}=-\mathbb{E}\left[h^{2}(w)\frac{\alpha z^{2}(\nu+1)}{2(\nu+z^{2})^{2}}\right],$
and Equation 16 can be simplified as
$\frac{\partial\ell(\alpha,\nu)}{\partial\nu}=\frac{1}{2}\log\left(\frac{\nu}{\nu+z^{2}}\right)+\frac{z^{2}-1}{2\left(\nu+z^{2}\right)}-h(w)\frac{\alpha z\sqrt{\nu+1}}{2\left(\nu+z^{2}\right)^{3/2}}+d_{\nu},$
where $d_{\nu}\coloneqq-\frac{1}{2}\psi\left(\frac{\nu}{2}\right)+\frac{1}{2}\psi\left(\frac{\nu+2}{2}\right)+2c_{\nu}*g_{1}(\nu)$. Hence, the Fisher information with respect to $\nu$ can be expressed as
$I_{\nu\nu}$ $=\mathbb{E}\left[\left(\frac{\partial\ell(\alpha,\nu)}{\partial\nu}\right)^{2}\right]$ (22)
$=\mathbb{E}\left[h^{2}(w)\frac{\alpha^{2}z^{2}(\nu+1)}{4\left(\nu+z^{2}\right)^{3}}\right]+\frac{1}{4}\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)^{2}\right]+\frac{1}{4}\mathbb{E}\left[\left(\frac{z^{2}-1}{\nu+z^{2}}\right)^{2}\right]+d^{2}_{\nu}$
$\qquad+\frac{1}{2}\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)\left(\frac{z^{2}-1}{\nu+z^{2}}\right)\right]+d_{\nu}\mathbb{E}\left[\log\frac{\nu}{\nu+z^{2}}\right]+d_{\nu}\mathbb{E}\left[\frac{z^{2}-1}{\nu+z^{2}}\right].$
After some integrations, we obtain
$\mathbb{E}\left[\frac{z^{2}-1}{\nu+z^{2}}\right]$ $=0$
$\mathbb{E}\left[\log\frac{\nu}{\nu+z^{2}}\right]$ $=\psi\left(\frac{\nu}{2}\right)-\psi\left(\frac{\nu+1}{2}\right)$
$\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)\left(\frac{z^{2}-1}{\nu+z^{2}}\right)\right]$ $=-\frac{1}{\nu^{2}+\nu}$
$\mathbb{E}\left[\left(\frac{z^{2}-1}{\nu+z^{2}}\right)^{2}\right]$ $=\frac{1}{\nu^{2}+3\nu}$
$\mathbb{E}\left[\left(\log\frac{\nu}{\nu+z^{2}}\right)^{2}\right]$ $=\left(\psi\left(\frac{\nu}{2}\right)-\psi\left(\frac{\nu+1}{2}\right)\right)^{2}+\psi^{(1)}\left(\frac{\nu}{2}\right)-\psi^{(1)}\left(\frac{\nu+1}{2}\right)$
$\mathbb{E}\left[h^{2}(w)\frac{\alpha^{2}z^{2}(\nu+1)}{4\left(\nu+z^{2}\right)^{3}}\right]$ $=\frac{\pi^{5/2}\Gamma\left(\frac{\nu}{2}+1\right)\left((\nu+3)H_{1}-(\nu+3)H_{2}-H_{3}+H_{4}\right)}{8\nu^{2}\Gamma\left(\frac{\nu+5}{2}\right)B\left(\frac{\nu}{2},\frac{1}{2}\right)B\left(\frac{\nu+1}{2},\frac{1}{2}\right)^{2}}$
$\mathbb{E}\left[h^{2}(w)\frac{\alpha z^{2}(\nu+1)}{2(\nu+z^{2})^{2}}\right]$ $=\frac{\pi\alpha\Gamma\left(\frac{\nu}{2}+1\right)^{2}\left((\nu+4)H_{5}-\frac{\left(\alpha^{2}(2\nu+3)+(\nu+3)\sigma_{\nu+1}^{2}\right)H_{6}}{\sigma_{\nu+1}^{2}}\right)}{8\left(\alpha^{2}+\sigma_{\nu+1}^{2}\right)\Gamma\left(\frac{\nu+1}{2}\right)\Gamma\left(\frac{\nu+5}{2}\right)},$
where
$H_{1}$ $=\,_{2}F_{1}\left(\frac{1}{2},\nu+1;\frac{\nu+3}{2};-\frac{\alpha^{2}}{\sigma_{\nu+1}^{2}}\right)$
$H_{2}$ $=\,_{2}F_{1}\left(\frac{1}{2},\nu+2;\frac{\nu+3}{2};-\frac{\alpha^{2}}{\sigma_{\nu+1}^{2}}\right)$
$H_{3}$ $=\,_{2}F_{1}\left(\frac{3}{2},\nu+1;\frac{\nu+5}{2};-\frac{\alpha^{2}}{\sigma_{\nu+1}^{2}}\right)$
$H_{4}$ $=\,_{2}F_{1}\left(\frac{3}{2},\nu+2;\frac{\nu+5}{2};-\frac{\alpha^{2}}{\sigma_{\nu+1}^{2}}\right)$
$H_{5}$ $=\,_{2}F_{1}\left(-\frac{1}{2},\nu+2;\frac{\nu+5}{2};-\frac{\alpha^{2}}{\sigma_{\nu+1}^{2}}\right)$
$H_{6}$ $=\,_{2}F_{1}\left(\frac{1}{2},\nu+2;\frac{\nu+5}{2};-\frac{\alpha^{2}}{\sigma_{\nu+1}^{2}}\right),$
which completes the proof. ∎
Goodness-of-fit for Skewed Student’s t Distribution
For simplicity, let $x_{1},x_{2},\ldots,x_{n}$ be the observations of $n$ control subjects. We denote $F(x|\alpha,\nu,\xi,\omega)$ as the cumulative function of this distribution where The parameters $\alpha,\nu,\xi,\omega$ are unknown. The definition of skewed Student’s $t$ distribution are defined in Section 2.2. To test the null hypothesis $H_{0}:x_{1},x_{2},\ldots,x_{n}$ are random samples from the skewed Student’s $t$ distribution, we propose the following testing process:
  1. (a)
    Estimate the parameters $\hat{\alpha},\hat{\nu},\hat{\xi},\hat{\omega}$ by maximising the likelihood function.
  2. (b)
    Compute the empirical distribution function $v_{i}=F(x|\hat{\alpha},\hat{\nu},\hat{\xi},\hat{\omega})$ for $i=1,2,\ldots,n$.
  3. (c)
    Compute $y_{i}=\Phi^{-1}(v_{i})$, where $\Phi$ is the standard normal cumulative distribution function.
  4. (d)
    Let $u_{i}=\Phi((y_{i}-\bar{y})/s_{y})$, where $\bar{y}=n^{-1}\sum_{i=1}^{n}y_{i}$ and $s_{y}^{2}=(n-1)^{-1}\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$. Let $u_{(1)},\ldots,u_{(n)}$ be the order statistics of $u_{1},\ldots,u_{n}$.
  5. (e)
    Calculate the Anderson statistic $A^{2}$ based on $u_{(1)},\ldots,u_{(n)}$, i.e.,
    $A^{2}:=-n-n^{-1}\sum_{i}^{n}[(2i-1)\log(u_{(i)})+(2n+1-2i)\log(1-u_{i})].$
    Modify $A^{2}$ into $A^{*}:=A^{2}(1+0.75/n+2.25/n^{2})$. If $A^{*}$ exceeds the critical value at a given significance level, reject $H_{0}$.
More details on the testing procedures and critical points can be found in Chen and Balakrishnan (1995). If we accept the null hypothesis $H_{0}$ at some significance level, then we can proceed with the comparison procedures between a single-subject and the control group described in the next section. If $H_{0}$ is rejected, one can employ a non-parametric method to compare a single-subject with the control group. However, the discussion of the non-parametric method is beyond the scope of this paper, we plan to extend this approach in future work.
Other results in Section 3.2
$n$ $z$ $t$ CG AD BIGPAST
FPR 50 0.0874 0.0788 0.0788 0.2214 0.0473
100 0.0677 0.0630 0.0630 0.1920 0.0389
200 0.0595 0.0571 0.0571 0.1972 0.0223
400 0.0533 0.0520 0.0520 0.1870 0.0144
TPR 50 0.9631 0.9574 0.9574 0.8090 0.9190
100 0.9820 0.9802 0.9802 0.8393 0.9571
200 0.9974 0.9968 0.9968 0.8586 0.9790
400 0.9983 0.9982 0.9982 0.8564 0.9819
ACC 50 0.9379 0.9393 0.9393 0.7938 0.9359
100 0.9572 0.9586 0.9586 0.8237 0.9591
200 0.9689 0.9698 0.9698 0.8307 0.9784
400 0.9725 0.9731 0.9731 0.8347 0.9837
Table 7: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under one-sided test (less). Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’ and ‘AD’ represent the Crawford-Garthwaite Bayesian and Anderson-Darling methods, respectively.
$n$ $z$ $t$ CG AD BIGPAST
FPR 50 0.0003 0.0002 0.0002 0.4511 0.0030
100 0.0000 0.0000 0.0000 0.4654 0.0044
200 0.0000 0.0000 0.0000 0.4699 0.0027
400 0.0000 0.0000 0.0000 0.4701 0.0021
TPR 50 0.6402 0.5907 0.5908 0.8229 0.8858
100 0.5445 0.5197 0.5198 0.7792 0.9206
200 0.5824 0.5693 0.5694 0.8491 0.9628
400 0.5389 0.5323 0.5323 0.8197 0.9670
ACC 50 0.8199 0.7953 0.7953 0.6859 0.9414
100 0.7722 0.7599 0.7599 0.6569 0.9581
200 0.7912 0.7847 0.7847 0.6896 0.9800
400 0.7694 0.7661 0.7661 0.6748 0.9825
Table 8: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under one-sided test(greater). Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’ and ‘AD’ represent the Crawford-Garthwaite Bayesian and Anderson-Darling methods, respectively.
$n$ $z$ $t$ CG AD BIGPAST
FPR 50 0.0467 0.0322 0.0322 0.0943 0.0313
100 0.0372 0.0286 0.0286 0.0523 0.0266
200 0.0324 0.0280 0.0280 0.0541 0.0231
400 0.0243 0.0220 0.0220 0.0392 0.0195
TPR 50 0.9353 0.9041 0.9041 0.9125 0.8854
100 0.9504 0.9358 0.9359 0.9302 0.9248
200 0.9617 0.9542 0.9541 0.9621 0.9450
400 0.9737 0.9700 0.9700 0.9568 0.9596
ACC 50 0.9443 0.9360 0.9359 0.9091 0.9270
100 0.9566 0.9536 0.9536 0.9389 0.9491
200 0.9647 0.9631 0.9631 0.9540 0.9610
400 0.9747 0.9740 0.9740 0.9588 0.9701
Table 9: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under two-sided test. The underlying density is normal density function. Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’ and ‘AD’ represent the Crawford-Garthwaite Bayesian and Anderson-Darling methods, respectively.
Other results in Section 3.3
(a) The density
(b) The type I error
(c) The power
(d) The accuracy
Figure 10: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=1$ and $\nu=10$.  Figure 10(a) shows the densities of skewed Student’s $t$ and normal distributions, the $\mu$ and $\sigma^{2}$ of normal distribution are equal to the mean and variance of skewed Student’s $t$ distribution. In fact, the normal distribution is the theoretical assumption of CG test when the sample comes from the skewed Student’s $t$ distribution. AD test is a nonparametric test; there is one result that is compared in the ‘two-sided’ category. TV distance is short for total variation distance. Each box plot summarises over 200 independent replications.
(a) The density
(b) The type I error
(c) The power
(d) The accuracy
Figure 11: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=1$ and $\nu=5$.  Figure 11(a) shows the densities of skewed Student’s $t$ and normal distributions, the $\mu$ and $\sigma^{2}$ of normal distribution are equal to the mean and variance of skewed Student’s $t$ distribution. In fact, the normal distribution is the theoretical assumption of CG test when the sample comes from the skewed Student’s $t$ distribution. AD test is a nonparametric test; there is one result that is compared in the ‘two-sided’ category. TV distance is short for total variation distance. Each box plot summarises over 200 independent replications.
(a) The density
(b) The type I error
(c) The power
(d) The accuracy
Figure 12: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=2$ and $\nu=5$.  Figure 12(a) shows the densities of skewed Student’s $t$ and normal distributions, the $\mu$ and $\sigma^{2}$ of normal distribution are equal to the mean and variance of skewed Student’s $t$ distribution. In fact, the normal distribution is the theoretical assumption of the CG test when the sample comes from the skewed Student’s $t$ distribution. AD test is a nonparametric test; there is one result that is compared in the ‘two-sided’ category. TV distance is short for total variation distance. Each box plot summarises over 200 independent replications.
Index label_name Index label_name
0 lh-bankssts 34 rh-bankssts
1 lh-caudalanteriorcingulate 35 rh-caudalanteriorcingulate
2 lh-caudalmiddlefrontal 36 rh-caudalmiddlefrontal
3 lh-cuneus 37 rh-cuneus
4 lh-entorhinal 38 rh-entorhinal
5 lh-frontalpole 39 rh-frontalpole
6 lh-fusiform 40 rh-fusiform
7 lh-inferiorparietal 41 rh-inferiorparietal
8 lh-inferiortemporal 42 rh-inferiortemporal
9 lh-insula 43 rh-insula
10 lh-isthmuscingulate 44 rh-isthmuscingulate
11 lh-lateraloccipital 45 rh-lateraloccipital
12 lh-lateralorbitofrontal 46 rh-lateralorbitofrontal
13 lh-lingual 47 rh-lingual
14 lh-medialorbitofrontal 48 rh-medialorbitofrontal
15 lh-middletemporal 49 rh-middletemporal
16 lh-paracentral 50 rh-paracentral
17 lh-parahippocampal 51 rh-parahippocampal
18 lh-parsopercularis 52 rh-parsopercularis
19 lh-parsorbitalis 53 rh-parsorbitalis
20 lh-parstriangularis 54 rh-parstriangularis
21 lh-pericalcarine 55 rh-pericalcarine
22 lh-postcentral 56 rh-postcentral
23 lh-posteriorcingulate 57 rh-posteriorcingulate
24 lh-precentral 58 rh-precentral
25 lh-precuneus 59 rh-precuneus
26 lh-rostralanteriorcingulate 60 rh-rostralanteriorcingulate
27 lh-rostralmiddlefrontal 61 rh-rostralmiddlefrontal
28 lh-superiorfrontal 62 rh-superiorfrontal
29 lh-superiorparietal 63 rh-superiorparietal
30 lh-superiortemporal 64 rh-superiortemporal
31 lh-supramarginal 65 rh-supramarginal
32 lh-temporalpole 66 rh-temporalpole
33 lh-transversetemporal 67 rh-transversetemporal
Table 10: The index of brain region of the Desikan-Killiany cortical atlas.